OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 4) with a slope of –1 in standard form.

OpenStudy (anonymous):

y - 4 = -1 ( x+1) y = -x-1+4 y = -x + 3

OpenStudy (anonymous):

–x + 4y = –1 x – 4y = –1 x + y = 3 x – y = –3 are what's there though

OpenStudy (anonymous):

my answer can be also wrote as x + y = 3